# A Novel Federated Learning-Based IDS for Enhancing UAVs Privacy and Security

### v2

@@ Abstract
Unmanned aerial vehicles (UAVs) operating within Flying Ad-hoc Networks (FANETs) encounter security challenges due to the dynamic and distributed nature of these networks. Previous studies predominantly focused on centralized intrusion detection, assuming a central entity responsible for storing and analyzing data from all devices. However, these approaches face challenges including computation and storage costs, along with a single point of failure risk, threatening data privacy and availability. The widespread dispersion of data across interconnected devices underscores the necessity for decentralized approaches. This paper introduces the Federated Learning-based Intrusion Detection System (FL-IDS), addressing challenges encountered by centralized systems in FANETs. FL-IDS reduces computation and storage costs for both clients and the central server, crucial for resource-constrained UAVs. Operating in a decentralized manner, FL-IDS enables UAVs to collaboratively train a global intrusion detection model without sharing raw data, thus avoiding the delay in decisions based on collected data, as is often the case with traditional methods. Experimental results demonstrate FL-IDS’s competitive performance with Central IDS (C-IDS) while mitigating privacy concerns, with the Bias Towards Specific Clients (BTSC) method further enhancing FL-IDS performance even at lower attacker ratios. Comparative analysis with traditional intrusion detection methods, including Local IDS (L-IDS), sheds light on FL-IDS’s strengths. This study significantly contributes to UAV security by introducing a privacy-aware, decentralized intrusion detection approach tailored to UAV networks. Moreover, by introducing a realistic dataset for FANETs and federated learning, our approach differs from others lacking high dynamism and 3D node movements or accurate federated data federations.
Index Terms:
@@
I Introduction
@@
Unmanned aerial vehicles (UAVs) deployed in flying ad-hoc networks (FANETs) [1] have attracted significant attention for their diverse applications, spanning surveillance, agriculture, disaster management, and communication infrastructure establishment [2, 3]. The dynamic and distributed nature of FANETs introduces substantial security challenges, necessitating robust intrusion detection systems [4, 5]. The increasing reliance on UAVs in critical applications underscores the unique security challenges posed by FANETs, emphasising the need for innovative intrusion detection solutions [6]. Unlike traditional networks, FANETs exhibit dynamic node movements, unreliable communication links, and a decentralised architecture [7].
@@
In this context, the Ad-Hoc On-Demand Distance Vector (AODV) protocol, a fundamental routing protocol in FANETs, faces challenges due to the mobility of UAVs, rendering traditional intrusion detection methods less effective [8]. The AODV protocol is susceptible to various routing attacks, including blackhole, sinkhole, and flooding attacks [9]. Traditional ML-based intrusion detection methods, such as Deep Neural Networks (DNN) and Convolutional Neural Networks (CNN), have been explored for their efficacy in identifying anomalous patterns [10, 11]. However, these methods often struggle with challenges related to communication costs, privacy concerns, and scalability.
@@
This paper introduces the Federated Learning-based Intrusion Detection System (FL-IDS), a novel intrusion detection approach tailored for FANETs, offering several key contributions to the field. Firstly, FL-IDS addresses privacy concerns by leveraging federated learning (FL) [12], a decentralised machine learning (ML) approach. This FL-based approach marks a departure from conventional intrusion detection paradigms. Secondly, our study contributes a realistic dataset based on an actual FANET characteristics, providing a more authentic foundation for analysis compared to synthetic datasets used in prior research.
@@
FL-IDS operates in a decentralised manner, enabling individual UAVs to train a global intrusion detection model collaboratively without sharing raw and sensitive data, addressing privacy concerns inherent in centralised approaches while harnessing collective intelligence. The architecture involves assigning default local models to each UAV, utilising client-specific data. After training local models, only the updated model weights are shared with a central server, which aggregates these weights to update the global model. This generated global model is responsible for detecting attacks, without delaying the decision for collecting data from clients as in traditional methods. This rapid decision-making capability is crucial in promptly addressing security threats, enhancing the effectiveness of our proposed intrusion detection system. Moreover, it considerably decreases communication cost, which is very important for FANETs which have frequent link breakages due to high nodes’ speeds. Last but not the least, it reduces the computation and storage cost both in clients and the central server which store, process the local data and aggregated data respectively. It is particularly important for resource-constrained UAVs running on small batteries. All these characterics make the proposed approach suitable for FANETs, while preserving data privacy.
@@
The novelty of FL-IDS lies in its effectiveness in detecting routing attacks, including sinkhole, blackhole, and flooding attacks, in the dynamic and decentralised FANET environment. Incorporating realistic network scenarios, considering 3D node movement and local data collection, enhances the applicability of FL-IDS. A comparative analysis with traditional intrusion detection methods, including Central IDS (C-IDS) and Local IDS (L-IDS), provides valuable insights into the strengths of FL-IDS.
@@
The remainder of this paper is organized as follows: In Section II, we delve into related works to contextualize our approach within existing literature. Section III provides background information on FANETs, routing attacks, and federated learning. Section IV details the proposed FL-IDS approach, explaining its architecture and functionality. Experimental results and analyses are presented in Section V. In Section VI, the proposed FL-IDS is discussed concerning its effectiveness, efficiency, communication cost, response time, privacy, and security. Additionally, a comparative analysis is presented with C-IDS and L-IDS across these specified criteria. Finally, Section VII concludes the paper, summarizing findings and outlining directions for future research.
II Related Work
@@
This section explores recent studies and approaches relevant to intrusion detection in FANETs. We explore latest advancements, methodologies, and datasets proposed in the literature, shedding light on how researchers have tackled the challenges of securing FANETs. From traditional ML to emerging FL-based solutions, we summarise the key insights from prior work. This understanding aids in evaluating the landscape and guides the development of our proposed intrusion detection approach.
@@
The use of ML (Support vector machines (SVM), Naive Bayes (NB), Linear Regression (LR) and Random Forest (RF)) for detecting DoS attacks against various components of a UAV is explored in [13]. The study utilised the DJI Phantom 4 drone dataset [14], which includes information collected from components such as GPS, gyro, flight controller, and battery module. While showing promising results of ML-based algorithms, the study also emphasized the need for reliable datasets to effectively train and validate ML models. Another study focusing on DoS attack detection is presented in [15]. This study employed the AWID2 dataset [16] for training three models (XGBoost, CatBoost, and LightGBM), with test conducted using real UAVs. The comparison of algorithms, considering factors like training time and area under the curve (AUC) metrics, revealed that LightGBM stands out among other models. Another study [17] also employs the traditional ML-based algorithms for detecting DDoS attacks and botnet activity in UAV networks.
@@
Another study [10] proposes a solution to detect GPS spoofing attacks categorized into two groups: static and dynamic. The dataset generated for both attacks collects GPS signal characteristics such as latitude, longitude, time from both authentic and spoofed experiments. The study employs traditional ML algorithms such as RF, K-Nearest Neighbors (KNN), SVM, Decision Trees (DT) and Neural Networks (NN), and uses the Sperman method for feature extraction. The results indicate that DT outperforms other algorithms, achieving a 92.36% detection rate and a 3.7% false positive rate. One of the main contributions of the study is the providing a real-time GPS spoofing detection solution that can be seamlessly integrated with standard receivers and ubiquitous modules, requiring no hardware modifications. A multi-agent-based IDS based on DT is proposed in [11]. A centrally located IDS collects data such as payload traffic, command and control traffic, and GPS data from UAVs. However, it has not been evaluated on a suitable dataset for FANETs. Hence, the authors leave the construction of a new dataset as future work.
@@
Many of the ML-based studies mentioned earlier rely on existing datasets, which may not be well-suited for FANETs. However, in two recent studies, researchers have addressed this gap in the literature by introducing datasets specifically tailored for FANETs. The first study [18] proposes an ML-based approach for detecting sybil attacks. To facilitate this, a dataset is constructed using the OMNET++ simulator [19], with a particular emphasis on capturing the 3D motion and density characteristics of FANETs. This dataset includes Received Signal Strength Difference (RSS) and Time Difference of Arrival (TDoA) derived from the physical layer, utilising ground monitoring stations. Experimental results show that this approach can detect sybil attacks with high accuracy (91%), with the false positive rate being less than 9% on average.
@@
Another attack dataset is introduced in [20]. This study is noteworthy as it is the first to address time delay attacks in FANETs, where delays are intentionally introduced in packet transmissions to the destination. The dataset comprises latency-related data collected from networks operating under normal conditions, simulated with the ONE simulator [21]. UAVs follow pre-planned routes determined by four different routing protocols. ML techniques are employed to identify anomalies in the dataset. The K-means clustering technique is subsequently applied to distinguish between malicious and benign nodes. The study demonstrates an accuracy of over 80% and less than 2.5% overhead in various network settings. However, this dataset is primarily designed for pre-planned flight paths, low speed (6 m/s) and 2D movement. These specifications may not align with the requirements of some other FANET applications, as the study is explicitly tailored for search and rescue missions.
@@
There are also studies that have explored the application of deep learning for intrusion detection in FANETs. It is shown in [22] that Convolutional Neural Networks (CNN) outperform traditional ML algorithms. The effectiveness of IDS was highlighted by an experimentally obtained accuracy of 99.50% and a prediction time of 2.77 ms on the UAV-IDS-2020 dataset [23]. In [24], Recurrent Neural Networks (RNN) are proposed to detect anomalies in the behavior of drones, such as unexpected changes in altitude, velocity, or trajectory. The model was trained using public datasets such as KDDCUP’99 [25] and NSL-KDD [26], which are not specific to FANETs. The model is deployed on each UAV and GBS, allowing attacks to be detected locally and then reported to the central IDS for verification. Experimental results show that the proposed method produces better results compared to LR and KNN.
@@
In [27], the Deep Reinforcement Learning (DRL) algorithm is employed to create an intrusion detection model using the CICIDS2017 dataset [28]. This model is deployed both on the central station and on each UAV. The approach introduces an offline learning system that automatically updates the model each time the UAV returns to its charging station, minimizing energy consumption. Another lightweight solution based on hierarchical-based SVM for detecting GPS and jamming attacks is proposed in [29]. The IDS on the UAVs uses a DL algorithm in conjunction with the Self-Taught Learning (STL) algorithm to extract significant features from movement logs. These logs contain routing information, velocity, and GPS location data collected during flights. Subsequently, SVM is employed for the detection process, and if an attack is detected, a Q-learning-based adaptive route learning algorithm is initiated on UAVs to return to a safe area. While the proposed approach claims to be well-suited for resource-constrained UAVs due to its utilization of DL, this claim lacks support from experimental results.
@@
Only a few studies [30, 31, 32] have applied federated learning to detect attacks in FANETs. In [30], the federated learning approach is employed to detect jamming attacks using two datasets. The first dataset, consisting of 3,000 samples and 8 features such as PDR, throughput, and received signal strength indicator (RSSI), is collected using Ns-3 [33]. The second dataset CRAWDAD [34], originally designed for VANETs, is modified to adopt a distributed data structure representing FANETs. Due to increased running time and communication costs when involving all clients in the training process, specific UAVs are selected for participation based on prioritization using the dumper-shaper technique. The accuracy obtained for the CRAWDAD dataset was approximately 82%, while for the FANET dataset, it reached around 89.5%. In contrast, traditional techniques yielded accuracies of 49.11% and 65.62% for the CRAWDAD and FANET datasets, respectively. In their extended study [31], a reinforcement federated learning-based technique is used to identify a defense strategy in unknown areas. This technique presents a different route to the desired location by avoiding the jamming attack region using the spatial retreat technique.
@@
In [32], an IDS is proposed for detecting flight anomalies and network attacks in UAV swarms. The Network Attack Detection subsystem employs a supervised approach, utilizing traditional algorithms (DT, RF, Gaussian Naive Bayes (NB), Multi Layer Perceptron (MLP), eXtreme Gradient Boosting (XGBoost), and LightGBM) to detect network attacks (blackhole, grayhole, and flooding). For the identification of flight anomalies, an unsupervised approach is adopted, leveraging federated learning to detect GPS jamming and spoofing attacks using the UAV Attack Dataset [35]. Various federated learning aggregation methods (FedAvg [36], FedAdagrad [37], FedAdam [37], and FedYogi [37]) are employed. Among them, FedAvg stands out for its robustness, achieving an F1-score of 0.887.
@@
There are also FL-based IDS proposed for IoT. A FL-based multiclass classifier is presented in [38]. The study examines three configurations as basic, balanced, and mixed, which are instances of splitting the ToN_IoT dataset [39] according to the IP addresses and types of attacks on IoT devices. The assessment takes into account the impact of aggregation functions such Fed+ [40] and FedAvg. The paper highlights the significance of an appropriate client/instance selection procedure to deal with issues in scenarios including non-independent, identically distributed, and highly skewed data. The evaluation’s findings emphasize the importance of this selection procedure and show that, in certain cases, using Fed+ significantly improves metrics over FedAvg.
@@
In the survey [41], an overview of studies based on FL in cyber security for various IoT applications, including the Internet of Vehicles, Internet of Drones, Industrial IoT, and the Internet of Healtcare Things is provided. The survey emphasizes the absence FL-based datasets in the literature, leading researchers to use and modify public datasets to mimic data federations. Even datasets collected from different environments [42] are employed for intrusion detection in IoT. However, it is imperative to utilize datasets that align with the specific characteristics of IoT or FANETs to develop more realistic intrusion detection solutions, as aimed in this study.
@@
There are also UAV-assisted applications where UAVs are being integrated into airspace operations for different ground-based entities and networks such as edge computing. However, the development of security strategies involving UAV assistance is still in its early stages and requires further research [43, 44].
@@
All related studies are summarized in Table I. As shown in the table, the majority of these studies train and evaluate their models on datasets not suitable for FANETs. Even in cases where new attack dataset for FANETs are created, their parameters are not very realistic due to simulating networks with a small number of nodes [30], 2D mobility models [20][29], or nodes with low speeds suitable for MANETs and VANETs but not FANETs [18][20]. Hence, such parameters may not accurately reflect the reality of FANETs. In contrast, our study simulates a total of 160 networks for three attacks (sinkhole, blackhole and flooding) with varying attacker ratios (from 5% to 25%) using 3D Gauss Markov Mobility (3D GM). To the best of our knowledge, this work represents the first proposal of a federated learning approach for these attacks. While a few federated learning-based studies have been proposed for jamming and spoofing attacks, their suitability has not been thoroughly evaluated in realistic network scenarios. Our study addresses this gap by exploring the use and suitability of federated learning for intrusion detection in FANETs, offering an extensive comparison with centralized and local IDS in terms of accuracy, communication cost, and privacy.
TABLE I: Outline of Related Studies
@@
Study Year Environment Dataset Attacks Method Pros (+) & Cons (-)
@@
[11] 2020 FANETs CICIDS2017 [28] Brute force, DoS, BotNet, Port Scanning ML + multi-agent-based IDS is proposed
SQL Injection, XSS, Heartbleed - the dataset is not suitable for FANETs
@@
[17] 2021 Cellular Connected UAV Networks CSE-CIC-IDS2018 [45] Brute force, DoS, DDoS ML + a study on cellular-based study
BotNet, Web Attack, Infiltration - the dataset is not suitable for FANETs and 5G networks
[13] 2022 UAVs DJI Phantom 4 [14] DoS attacks ML + different types of ML-based algorithms are evaluated
- focused on sensor-based attacks rather than FANETs-specific attacks
[15] 2022 UAVs AWID2 [16] (for training and testing in simulation) DoS attacks ML + various ML-based algorithms are assessed with real UAVs
individual UAV data transmission (for real testing) + providing hyperparameter optimization using Bayesian algorithm

### v3

@@ Abstract
Unmanned aerial vehicles (UAVs) operating within Flying Ad-hoc Networks (FANETs) encounter security challenges due to the dynamic and distributed nature of these networks. Previous studies focused predominantly on centralized intrusion detection, assuming a central entity responsible for storing and analyzing data from all devices. However, these approaches face challenges including computation and storage costs, along with a single point of failure risk, threatening data privacy and availability. The widespread dispersion of data across interconnected devices underscores the need for decentralized approaches. This paper introduces the Federated Learning-based Intrusion Detection System (FL-IDS), addressing challenges encountered by centralized systems in FANETs. FL-IDS reduces computation and storage costs for both clients and the central server, which is crucial for resource-constrained UAVs. Operating in a decentralized manner, FL-IDS enables UAVs to collaboratively train a global intrusion detection model without sharing raw data, thus avoiding delay in decisions based on collected data, as is often the case with traditional methods. Experimental results demonstrate FL-IDS’s competitive performance with Central IDS (C-IDS) while mitigating privacy concerns, with the Bias Towards Specific Clients (BTSC) method further enhancing FL-IDS performance even at lower attacker ratios. Comparative analysis with traditional intrusion detection methods, including Local IDS (L-IDS), sheds light on the strengths of FL-IDS. This study significantly contributes to UAV security by introducing a privacy-aware, decentralized intrusion detection approach tailored to UAV networks. Moreover, by introducing a realistic dataset for FANETs and federated learning, our approach differs from others lacking high dynamism and 3D node movements or accurate federated data federations.
keywords:
@@
journal: Internet of Things; Engineering Cyber Physical Human Systems
[…]
[label3]organization=Department of Computer Science, University of York, city=York, postcode=YO10 5GH, country=United Kingdom
1 Introduction
@@
Unmanned aerial vehicles (UAVs) deployed in flying ad hoc networks (FANETs) [1] have attracted significant attention for their diverse applications, spanning surveillance, agriculture, disaster management and the establishment of communication infrastructures [2, 3]. The dynamic and distributed nature of FANETs introduces substantial security challenges, which require robust intrusion detection systems [4, 5]. The increasing reliance on UAVs in critical applications underscores the unique security challenges posed by FANETs, emphasizing the need for innovative intrusion detection solutions [6]. Unlike traditional networks, FANETs exhibit dynamic node movements, unreliable communication links, and a decentralized architecture [7].
@@
In this context, the Ad-Hoc On-Demand Distance Vector (AODV) protocol, a fundamental routing protocol in FANETs, faces challenges due to the mobility of UAVs, rendering traditional intrusion detection methods less effective [8]. The AODV protocol is susceptible to various routing attacks, including blackhole, sinkhole, and flooding attacks [9]. Traditional ML-based intrusion detection methods, such as Deep Neural Networks (DNN) and Convolutional Neural Networks (CNN), have been explored for their efficacy in identifying anomalous patterns [10, 11]. However, these methods often face challenges related to communication costs, privacy concerns, and scalability.
@@
This paper introduces the Federated Learning-based Intrusion Detection System (FL-IDS), a novel intrusion detection approach tailored for FANETs, offering several key contributions to the field. Firstly, FL-IDS addresses privacy concerns by leveraging federated learning (FL) [12], a decentralized machine learning (ML) approach. This FL-based approach marks a departure from conventional intrusion detection paradigms. Secondly, our study contributes a realistic dataset based on actual FANET characteristics, providing a more authentic foundation for analysis compared to synthetic datasets used in previous research.
@@
FL-IDS operates in a decentralized manner, enabling individual UAVs to train a global intrusion detection model collaboratively without sharing raw and sensitive data, addressing privacy concerns inherent in centralized approaches while harnessing collective intelligence. The architecture involves assigning default local models to each UAV, utilising client-specific data. After training local models, only the updated model weights are shared with a central server, which aggregates these weights to update the global model. This generated global model is responsible for detecting attacks without delaying the decisions by collecting raw data from clients, as in traditional methods. This rapid decision-making capability is crucial in promptly addressing security threats, enhancing the effectiveness of our proposed intrusion detection system. Moreover, it considerably decreases communication cost, which is very important for FANETs which have frequent link breakages due to high nodes’ speeds.
The role of a central server remains crucial in FANET architectures, complementing the decentralized nature of FL-IDS. While FANETs often utilize peer-to-peer connections for coordination and collaboration among UAVs, they also frequently incorporate a central server, especially for tasks like data collection and relaying information to a command control center [1]. It is both practical and beneficial for certain FANET configurations due to its role in aggregating model updates rather than handling raw data transmissions. The central server is typically implemented at a Ground Base Station (GBS) or a stable UAV acting as a cluster head, as these entities possess greater computational power, stable energy supplies, and access to reliable communication infrastructure, making them well-suited for aggregating model updates. The use of a central server also provides critical benefits for resource-constrained UAVs. By offloading the task of aggregating model updates to the server, the computational and storage demands on individual UAVs are reduced. This is especially vital in FANETs, where UAVs often operate on limited battery power and computational resources.
@@
The novelty of FL-IDS lies in its effectiveness in detecting routing attacks, including sinkhole, blackhole, and flooding attacks, in the dynamic and decentralised FANET environment. Additionally, the Bias Towards Specific Clients (BTSC) method, which adjusts the model’s focus to prioritize clients with superior attack detection capabilities, further enhances FL-IDS’s performance, even in scenarios with a low attacker ratio. Incorporating realistic network scenarios, considering 3D node movement and local data collection, enhances the applicability of FL-IDS. A comparative analysis with traditional intrusion detection methods, including Central IDS (C-IDS) and Local IDS (L-IDS), provides valuable insights into the strengths of FL-IDS.
The main contributions of our study are summarized as follows:
  • 1.
    We propose a novel FL-IDS specifically designed for FANETs, addressing privacy concerns by enabling collaborative learning without sharing raw data. The decentralized operation reduces communication costs, making it suitable for the highly dynamic and resource-constrained nature of FANETs.
  • 2.
    We introduce a realistic dataset based on actual FANET characteristics, incorporating 3D node movement and local data collection. This improves the system’s applicability to real-world environments and includes critical routing attacks such as sinkhole, blackhole, and flooding attacks.
  • 3.
    The proposed system demonstrates effectiveness in detecting critical routing attacks—including sinkhole, blackhole, and flooding—while the Bias Towards Specific Clients (BTSC) method enhances FL-IDS performance even at scenarios with lower attacker ratios.
  • 4.
    A comprehensive comparative analysis with traditional IDS approaches, including Central IDS (C-IDS) and Local IDS (L-IDS), highlights the advantages of FL-IDS in effectiveness, efficiency, and privacy preservation.
@@
The remainder of this paper is organized as follows. In Section 2, we delve into related works to contextualize our approach within existing literature. Section 3 provides background information on FANETs, routing attacks, and federated learning. Section 4 details the proposed FL-IDS approach, explaining its architecture and functionality. The experimental results and analyses are presented in Section 5. In Section 6, the proposed FL-IDS is discussed concerning its effectiveness, efficiency, communication cost, response time, privacy, and security. In addition, a comparative analysis is presented with C-IDS and L-IDS across these specified criteria. Finally, Section 7 concludes the paper, summarizing findings and outlining directions for future research.
2 Related Work
@@
This section explores recent studies and approaches relevant to intrusion detection in FANETs. We explore the latest advances, methodologies, and datasets proposed in the literature, shedding light on how researchers have tackled the challenges of FANET security. From traditional ML to emerging FL-based solutions, we summarize the key insights from prior work. This understanding helps to evaluate the landscape and guides the development of our proposed intrusion detection approach.
@@
The use of ML (Support Vector Machines (SVM), Naive Bayes (NB), Linear Regression (LR), and Random Forest (RF)) for detecting DoS attacks against various components of a UAV is explored in [13]. The study used the DJI Phantom 4 drone dataset [14], which includes information collected from components such as GPS, gyro, flight controller, and battery module. While showing promising results for ML-based algorithms, with the best accuracy of 97.84% achieved by RF algorithm, the study also emphasized the need for reliable datasets to effectively train and validate ML models. Another study focusing on DoS attack detection is presented in [15]. This study employed the AWID2 dataset [16] for training three models (XGBoost, CatBoost, and LightGBM), with tests conducted using real UAVs. The comparison of algorithms, considering factors like training time and area under the curve (AUC) metrics, revealed that LightGBM stands out among other models, achieving an AUC of 99.83%. Another study [17] also employs traditional ML-based algorithms for detecting DDoS attacks and botnet activity in UAV networks. Among these ML methods, the Decision Tree (DT) algorithm demonstrated exceptional performance, achieving a maximum accuracy of 99.99%.
@@
Another study [10] proposes a solution to detect GPS spoofing attacks categorized into two groups: static and dynamic. The dataset generated for both attacks collects GPS signal characteristics such as latitude, longitude, and time from both authentic and spoofed experiments. The study employs traditional ML algorithms such as RF, K-Nearest Neighbors (KNN), SVM, DT, and Neural Networks (NN) and uses the Spearman method for feature extraction. The results indicate that DT outperforms other algorithms, achieving a 92.36% detection rate and a 3.7% false positive rate. One of the main contributions of the study is providing a real-time GPS spoofing detection solution that can be seamlessly integrated with standard receivers and ubiquitous modules, requiring no hardware modifications. A multi-agent-based IDS based on DT is proposed in [11]. A centrally located IDS collects data such as payload traffic, command and control traffic, and GPS data from UAVs. The study reported achieving an accuracy of 100%; however, it has not been evaluated on a suitable dataset for FANETs. Hence, the authors leave the construction of a new dataset as future work.
@@
Many of the ML-based studies mentioned earlier rely on existing datasets, which may not be well-suited for FANETs. However, in two recent studies, researchers have addressed this gap in the literature by introducing datasets specifically tailored for FANETs. The first study [18] proposes an ML-based approach for detecting Sybil attacks. To facilitate this, a dataset is constructed using the OMNET++ simulator [19], with a particular emphasis on capturing the 3D motion and density characteristics of FANETs. This dataset includes Received Signal Strength Difference (RSS) and Time Difference of Arrival (TDoA) derived from the physical layer, utilizing ground monitoring stations. Experimental results show that this approach can detect Sybil attacks with high accuracy (91%), with the false positive rate being less than 9% on average.
@@
Another attack dataset is introduced in [20]. This study is noteworthy as it is the first to address time delay attacks in FANETs, where delays are intentionally introduced in packet transmissions to the destination. The dataset comprises latency-related data collected from networks operating under normal conditions, simulated with the ONE simulator [21]. UAVs follow pre-planned routes determined by four different routing protocols. ML techniques are employed to identify anomalies in the dataset. The K-means clustering technique is subsequently applied to distinguish between malicious and benign nodes. The study demonstrates an accuracy of over 80% and less than 2.5% overhead in various network settings. However, this dataset is primarily designed for pre-planned flight paths, low speed (6 m/s), and 2D movement. These specifications may not align with the requirements of some other FANET applications, as the study is explicitly tailored for search and rescue missions.
@@
There are also studies that have explored the application of deep learning for intrusion detection in FANETs. It is shown in [22] that Convolutional Neural Networks (CNNs) outperform traditional ML algorithms. The effectiveness of IDS was highlighted by an experimentally obtained accuracy of 99.50% and a prediction time of 2.77 ms on the UAV-IDS-2020 dataset [23]. In [24], Recurrent Neural Networks (RNNs) are proposed to detect anomalies in the behavior of drones, such as unexpected changes in altitude, velocity, or trajectory. The model was trained using public datasets such as KDDCup’99 [25] and NSL-KDD [26]. Although it achieved an average accuracy of 95%, these datasets are not specifically tailored to FANET environments. The model is deployed on each UAV and GBS, allowing attacks to be detected locally and then reported to the central IDS for verification. The experimental results show that the proposed method produces better results compared to LR and KNN.
In [27], the Deep Reinforcement Learning (DRL) algorithm is employed to create an intrusion detection model using the CICIDS2017 dataset [28], achieving an accuracy of 99.70%. This model is deployed both on the central station and on each UAV. The approach introduces an offline learning system that automatically updates the model each time the UAV returns to its charging station, minimizing energy consumption.
@@
Another lightweight solution based on hierarchical SVM for detecting GPS and jamming attacks is proposed in [29]. The IDS on the UAVs uses a DL algorithm in conjunction with the Self-Taught Learning (STL) algorithm to extract significant features from movement logs. These logs contain routing information, velocity, and GPS location data collected during flights. While SVM achieved an accuracy of over 72%, when combined with STL, the accuracy improves significantly, reaching over 92%. If an attack is detected, a Q-learning-based adaptive route learning algorithm is initiated on UAVs to return to a safe area. While the proposed approach claims to be well-suited for resource-constrained UAVs due to its utilization of DL, this claim lacks support from experimental results.
@@
Only a few studies [30, 31, 32] have applied federated learning to detect attacks in FANETs. In [30], the federated learning approach is employed to detect jamming attacks using two datasets. The first dataset, consisting of 3,000 samples and 8 features such as PDR, throughput, and received signal strength indicator (RSSI), is collected using Ns-3 [33]. The second dataset CRAWDAD [34], originally designed for VANETs, is modified to adopt a distributed data structure that represents FANETs. Due to increased running time and communication costs when involving all clients in the training process, specific UAVs are selected for participation based on prioritization using the dumper-shaper technique. The accuracy obtained for the CRAWDAD dataset was approximately 82%, while for the FANET dataset, it reached around 89.5%. In contrast, traditional techniques yielded accuracies of 49.11% and 65.62% for the CRAWDAD and FANET datasets, respectively. In their extended study [31], a reinforcement federated learning-based technique is used to identify a defense strategy in unknown areas. This technique presents a different route to the desired location by avoiding the jamming attack region using the spatial retreat technique.
@@
In [32], an IDS is proposed to detect flight anomalies and network attacks in UAV swarms. The Network Attack Detection subsystem employs a supervised approach, utilizing traditional algorithms (DT, RF, Gaussian Naive Bayes (NB), Multi-Layer Perceptron (MLP), eXtreme Gradient Boosting (XGBoost), and LightGBM) to detect network attacks (blackhole, grayhole, and flooding). For the identification of flight anomalies, an unsupervised approach is adopted, leveraging federated learning to detect GPS jamming and spoofing attacks using the UAV Attack Dataset [35]. Various federated learning aggregation methods (FedAvg [36], FedAdagrad [37], FedAdam [37], and FedYogi [37]) are employed. Among them, FedYogi stands out for its robustness, achieving an F1-score of 0.904.
@@
FL-based IDSs have also been proposed for IoT. An FL-based multiclass classifier is presented in [38]. The study examines three configurations as basic, balanced, and mixed, which are instances of splitting the ToN_IoT dataset [39] according to IP addresses and types of attacks on IoT devices. The assessment takes into account the impact of aggregation functions such as Fed+ [40] and FedAvg. The paper highlights the significance of an appropriate client/instance selection procedure to deal with issues in scenarios including non-independent, identically distributed, and highly skewed data. The evaluation findings emphasize the importance of this selection procedure and show that, in certain cases, using Fed + significantly improves the metrics over FedAvg, achieving an accuracy of over 80%. Another study [41] presented a secure and efficient FL approach that incorporates blockchain technology to enhance security, privacy, and trust in IDSs for IoT networks. It introduces a secure aggregation algorithm using Secure Multi-Party Computation (SMPC), a cryptographic technique that enables multiple parties to collaboratively compute a function on their private data without revealing any individual inputs. In addition, a blockchain-based reputation mechanism is used to ensure data privacy and model integrity. The proposed approach achieved an accuracy of 99%. In another study [42], blockchain and edge computing were applied in V2X networks to ensure node trustworthiness and reduce communication delays.
@@
In the survey [43], an overview of FL-based cybersecurity studies is provided for various IoT applications, including the Internet of Vehicles, the Internet of Drones, Industrial IoT, and the Internet of Healtcare Things. The survey emphasizes the absence of FL-based datasets in the literature, which leads researchers to use and modify public datasets to mimic data federations. Even datasets collected from different environments [44] are used for intrusion detection in IoT. However, it is imperative to utilize datasets that align with the specific characteristics of IoT or FANETs to develop more realistic intrusion detection solutions, as aimed in this study.
@@
There are also UAV-assisted applications where UAVs are being integrated into airspace operations for different ground-based entities and networks, such as edge computing. However, the development of security strategies that involve the assistance of UAVs is still in its early stages and requires further research [45, 46].
@@
All related studies are summarized in Table 1. As shown in the table, the majority of these studies train and evaluate their models on datasets not suitable for FANETs. Even in cases where new attack datasets for FANETs are created, their parameters are not very realistic due to simulating networks with a small number of nodes [30], 2D mobility models [20][29], or nodes with low speeds suitable for MANETs and VANETs but not FANETs [18][20]. In contrast, FANET datasets possess distinct characteristics that set them apart from MANET and VANET datasets. These include high-speed node mobility in 3D space, frequent topology changes, and variable link reliability due to UAV dynamics. Such factors introduce significant differences in network behavior and attack patterns compared to MANETs and VANETs, which typically assume simpler 2D mobility models and more stable link conditions. Consequently, datasets designed for FANETs must incorporate realistic 3D mobility models, account for high-speed UAV nodes, and consider frequent link disruptions to ensure the reliable evaluation of IDSs in FANET environments. Our study simulates a total of 160 networks for three attacks (sinkhole, blackhole, and flooding) with varying attacker ratios (from 5% to 25%) using 3D Gauss Markov Mobility (3D GM). To the best of our knowledge, this work represents the first proposal of a federated learning approach for these attacks. Although some federated learning-based studies have been proposed for jamming and spoofing attacks, their suitability has not been thoroughly evaluated in realistic network scenarios. Our study addresses this gap by exploring the use and suitability of federated learning for intrusion detection in FANETs, offering an extensive comparison with centralized and local IDS in terms of accuracy, communication cost, and privacy.
Table 1: Outline of Related Studies
@@
Study Year Environment Dataset Attacks Method Perf. Metrics Pros (+) & Cons (-)
@@
[11] 2020 FANETs CICIDS2017 [28] Brute force, DoS, BotNet, Port Scanning ML Accuracy +Multi-agent-based IDS is proposed
SQL Injection, XSS, Heartbleed -Dataset is not suitable for FANETs
@@
[17] 2021 Cellular Connected UAV Networks CSE-CIC-IDS2018 [47] Brute force, DoS, DDoS ML Accuracy, Precision +A study on cellular-based study
BotNet, Web Attack, Infiltration Recall, F1-score, FNR -The dataset is not suitable for FANETs and 5G networks
[13] 2022 UAVs DJI Phantom 4 [14] DoS attacks ML Accuracy, Precision +Evaluates various ML algorithms
Recall, AUC -Focused on sensor-based attacks rather than FANETs-specific attacks
[15] 2022 UAVs AWID2 [16] (for training and testing in simulation) DoS attacks ML Accuracy, Precision +Various ML-based algorithms are assessed with real UAVs
individual UAV data transmission (for real testing) Recall, F1-score, FPR +Providing hyperparameter optimization using Bayesian algorithm